OpenStudy (anonymous):

Three consecutive odd integers have a sum of 39. Find the integers.

OpenStudy (anonymous):

Use the same equation as before 3x+3 = 39 it applies to all problems that are like this

OpenStudy (anonymous):

but the integers have to be odd

OpenStudy (anonymous):

Oh yes they do...

OpenStudy (anonymous):

Then the equation is (x) + (x+2) + (x+4) =39 or 3x+6 = 39

OpenStudy (anonymous):

I got 1,3, and 5 and the answer was wrong :(

OpenStudy (anonymous):

Well 1 3 and 5 don't add up to 39

OpenStudy (anonymous):

they add up to 9

OpenStudy (anonymous):

i solved it wrong woops

OpenStudy (anonymous):

Haha ya

OpenStudy (anonymous):

i got it now thanks !

OpenStudy (anonymous):

Yup. Anything else?

OpenStudy (anonymous):

Medal maybe?